Presidential Review 2003

Sky Weber

February 26, 2003

‘If there is one underlying principle that must guide the next twelve months it is to remember that our members are not customers’. Some of you were here last year and may remember Richard Bray; the previous President ending last year’s Presidential Review on that somewhat humble request. Although many people disagreed with it, I feel, for reasons that will become apparent, that it is a very fitting way to begin this year’s review.

One of the last tasks for a President before they hand over is to present a review of the Society’s activities over the past year. I now stand before you to talk about where we have showed strength and camaraderie and where we have not been able to rise to the challenges that have been set before us. I’ll try and keep it as short as possible as I’m sure you all have much better things to do then listen to me here, like listen to me present the Film Quiz over at the bar in a few moments time. On a slightly more serious note, I would like to keep the formality that is the Presidential Review as short as possible to ensure that you all have the opportunity to question me and the rest of the exec about anything you would like to know about the Society and our stint in office.

This year started for us with an extra-ordinary event, the special screening of One Giant Leap. This screening, attended by the directors, provided me with a very step learning curve into the logistics of running a society as large as ours, the politics within it and the external forces that ultimately govern it. Being President has given me the privilege of working with some fantastic people, who may at times have given me a headache, but it has been an experience that I shall never forget, and am sure will draw upon for the rest of my life. Indeed there have been some fantastic achievements within the society this past year that we should be proud of.

Technical improvements, and impairments have been carried out this year under the very watchful and occasionally bloodshot eye of Amanda Window. The biggest project the technical team undertook this year has been the upgrade to the sound system. This has involved new equipment from one end of the sound chain, with new surround speakers and amplifiers, right through to the source itself with the soon to be commissioned Sony Dynamic Digital Sound. SDDS is the 3^{rd} major digital sound format and Warwick Student Cinema will be the only Student Cinema in the country to boast this superior format and one of only a handful in the region, reaffirming our status as one of the technical leaders in our field.

Another area that has seen exponential growth has been the Society’s IT infrastructure. The IT team headed up by the demure JP has evolved the limited IT resources of the society from a single account on the University network to a complete in-house network serving all the society’s web hosting, e-mail, and database requirements. Although the computers used to provide these services are very old, slow and decrepit, when networked together they can accomplish tasks greater then collective abilities put together.

I know this concept of synergy is almost a cliché but it is very true and one which we must not forget. There will always be times when a cog is unable to perform its role to the standard that has been set by their peers, and more importantly by themselves. We must remember that when we are part of a bigger society, we should not loose sight of the entire picture. We must come together and help each other even if it means that we must make sacrifices ourselves. However, for this concept to work it is imperative that we admit when we are wrong and have made mistakes. There is no one in this room that could say, with their hand on their heart, that they are perfect and have never done anything they regret or anything that they feel they could have done better. We all make mistakes, we must move on and learn from them and admit when we are wrong.

Socials as usual have provided a very important aspect to our society. This is the biggest society on campus and it is easy to get lost within it and it can be very difficult to break into already formed cliques. To ensure that we can get people more involved we must make this feel more like a society. With the new IT system with have seen the introduction of the ability to mass mail our members. This has made it even easier to keep them up to date with the goings on within the society. It also makes informing them about special events such as The AllNighter and preview screening less troublesome then under the traditional paper methods of the past. However, despite the clear benefits that IT gives us by bringing us closer, we must be cautious not to let it drag us apart. One area where this could be a potential issue concerns the fortnightly General Meetings where the advent of IT has the power to lead to their ultimate demise. The introduction of on-line rotas and and an improved web page have the potential to remove some of the main reasons for coming to these meetings. However, the General Meetings should be more then rotas and news, they should provide a forum for everyone to meet face to face without the usual restrictions that a computer can imposes on us.

So. Going back to the point I started my review with: Do we have members or customers? The easiest way to answer that question is to ask ourselves, are we a society or are we a service? I think to be able to answer that question correctly we should break the society into two distinct categories. Number 1: The Associates. You guys and gals make the society what it is, it is with your dedication and commitment that we are able to show films almost every night of the week and we need to repay that dedication by giving you more then just something to put on your CV.

It is very easy to get carried away with being highly involved in the society, it is very important to enjoy yourself every once in a while and to have fun. This term has seen film showing times put back by half an hour to allow the Associates working shows to have more time afterwards to socialise and add more fun to being a part of our family. So on this level I can positivly say that we are a society.

The second group of people in our society is populated by The Members. These are the people who purely come to watch films and enjoy themselves. Some of you here, and indeed most of those over in The Union feel that this group of people are merely customers and not members. Why do they feel this? Because to them all we do is provide them with a cinema. On a level this may be true, but we must remember that they pay for this service, not only in terms of buying a ticket but also by becoming a member. This year the Union imposed a ‘SocsFed’ tax on joining societies, raising the cost of joining a society from £2.50 to almost £8 and we must strive to add as much value as possible to memberships now more then ever in order to counteract the decrease in membership that we, along with most societies have experienced this year. This can be done by introducing more speaker evenings, and by arranging more preview screenings both of which have been very scarce this year due to forces out of our control. We can also work with other societies to give our members a bigger package that consists of more then just a ticket to see a film. When we give up the fight and admit to The Union that we are just a money making service and our members are in fact just customers, as they would like us to, we shall divorce ourselves from this very important part of our family. This part of the society gives us meaning; it provides us with the reason for the societies existence. Film is a very powerful media and has the capability to affect people in many ways. It has the power to make us laugh, make us cry and most importantly open our minds and make us think.

On an affirming crescendo Warwick Student Cinema is a magnificent society and I mean that in the most positive way, true to our society’s embracing tradition.

Thank you